# Residual Diffusion Modeling for Km-scale Atmospheric Downscaling

### v3

@@ 5.2.1 Regression on the mean
In order to predict the conditional mean ${\boldsymbol{\mu}}=\mathbbm{E}[{\bf x}|{\bf y}]$, we resort to UNet-regression. A UNet network is supervised with training data $\{({\bf x}_{n},{\bf y}_{n})\}_{n=1}^{N}$ to learn the regression. We adopt a UNet architecture that is commonly used in denoising diffusion models. This particular UNet [53] incorporates attention layers and residual layers, allowing it to effectively capture both short and long-range dependencies in the data see section 2 and S1. Mean-Squared-Error (MSE) loss is optimized for training.
@@ 5.3.1 Dataset
The input (conditioning) dataset is taken from ERA5 reanalysis [25], a global dataset at spatial resolution of about 25km and a temporal resolution of 1h, the latter matches the target data listed below. To facilitate training, we interpolate the input data onto the curvirectangular grid of CWA with bilinear interpolation (with a rate of 4x), which results in $36\times 36$ pixels over the region of Taiwan. Each sample in the input dataset consists of 20 channels of information see S2. This includes four pressure levels (500 hPa, 700 hPa, 850 hPa, and 925 hPa) with four corresponding variables: temperature, eastward and northward components horizontal wind vector, and Geopotential Height. Additionally, the dataset includes single level variables such as 2 meter Temperature, 10 meter wind vector, total column water vapor.
@@
The target dataset used in this study is a subset of the proprietary RWRF model data (Radar Data Assimilation with WRFDA  [^1] https://www.mmm.ucar.edu/models/wrfda). The RWRF model is one of the operational regional Numerical Weather Prediction (NWP) models developed by CWA [13], which focuses on radar Data Assimilation (DA) in the vicinity of Taiwan. Assimilating radar data is a common strategy used in regional weather prediction, which helps constrain especially stochastic convective processes such as mesoscale convective systems and short-lived thunderstorms. The WRF - CWA system uses a nested 2km domain in a 10km that is driven by a global model (GFS) as boundary conditions [13]. By incorporating radar data [12], RWRF improves the short-term prediction of high-impact weather events. The radar observations possess high spatial resolution of approximately 1km and temporal resolutions of 5-10 minutes at a convective scale. These observations provide wind information (radial velocity) as well as hydrometers (radar reflectivity) usefully, with a particular emphasis on the lower atmosphere. The radar data assimilation relies on the availability of reliable and precise observations, which contributes significantly to enhance the accuracy and performance of the applied deep learning algorithms in the context of NWP applications.
@@
The target dataset covers a duration of 52 months, specifically from January 2018 to April 2022. It has a temporal frequency of one hour and a spatial resolution of 2km. We use only the inner (nested) 2km domain, which has $448\times 448$ pixels, projected using the Lambert conformal conical projection method around Taiwan. The geographical extent of the dataset spans from approximately 116.371°E to 125.568°E in longitude and 19.5483°N to 27.8446°N in latitude. We sub-selected 4 channels (variables) as the target variables, that are most relevant for practical forecasting: temperature at 2 meter above the surface, the horizontal winds at 10 meter above the surface and the 1h maximum radar reflectivity - a surrogate of expect precipitation. Notably, the radar reflectivity channel is not present in the input data and needs to be predicted based on the other channels, making its prediction strictly generative. The radar reflectivity data also exhibits a distinct distribution compared to the other output channels, with positive values and a prominent zero-mode consistent with typical non-raining conditions.
@@
Initially, the target data is provided in the NetCDF format, which is the output of the WRFDA assimilation process. Subsequently, vertical interpolated from sigma levels to isobaric levels then save as a custom CWA DMS format. As part of the preprocessing steps, the data is converted to the Hadoop Distributed File System (HDFS) format. Additionally, any missing or corrupted data points represented by "inf" or "nan" values are eliminated from the dataset. This leads to a reduction in the number of samples from 37,944 to 33,813.
@@
To avoid over-fitting we divide the data into training and testing sets. Three years of data 2018-2020 are used for training (2,4154 samples total). For testing we use the full fourth year, 2021, as well as the first four months (January to April) of 2022.
@@ 5.3.2 Network architecture and training
The CorrDiff method has two step learning approach. To ensure compatibility and consistency, we employ the same UNet architecture used in EDM (Elucidated Diffusion Models) for both the regression and diffusion networks(see S1). This architecture is based on the UNet model proposed in [53]. We enhance the UNet by increasing its size to include 6 encoder layers and 6 decoder layers. The base embedding size is set to 128, and it is multiplied over channels according to the pattern [1,2,2,2,2]. The attention resolution is defined as 28. For time embedding in the diffusion process, we utilize Fourier-based position embedding. However, in the regression network, time embedding is disabled. No data augmentation techniques are employed during training. Overall, the UNet architecture comprises 80 million parameters. Additionally, we introduce 4 channels for sinusoidal positional embedding.
During the training phase, we use the Adam optimizer with a learning rate of $2\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.99$. Exponential moving averages (EMA) with a rate of $\eta=0.5$ are applied, and dropout with a rate of 0.13 is utilized. The regression network solely receives the input conditioning channels. On the other hand, the diffusion training incorporates the 20 input conditioning channels from the coarse-resolution ERA5 data, which are concatenated with 4 noise channels to generate the output for each denoiser. For diffusion training, we adopt the Elucidated Diffusion Model (EDM), a continuous-time diffusion model. During training, EDM randomly selects the noise variance such that $\ln(\sigma(t))\sim\mathcal{N}(0,1.2^{2})$ and aims to denoise the samples per mini-batch. EDM is trained for 100 million steps, while the regression network is trained for 30 million steps. The training process is distributed across 16 DGX nodes, each equipped with 8 A100 GPUs, utilizing data parallelism and a total batch size of 512. The total training time for regression and diffusion models was 7 days that amounts to approximately $21,504$ GPU-hours.
For sampling purposes, we employ the second-order stochastic sampler provided by EDM. This sampler performs 18 steps, starting from a maximum noise variance of $\sigma_{\max}=800$ and gradually decreasing it to a minimum noise variance of $\sigma_{\min}=0.002$. We adopt the rest of hyperparamaters from EDM as listed in [28].
@@ 5.4 Evaluation criterion
Probabilistic predictions aim to maximize sharpness subject to calibration [44]. Qualitatively, calibration means that the likelihood of observing the true value is the same as observing a member drawn from the ensemble. A necessary condition for calibration is that the spread-error relationship be 1-to-1 when averaged over sufficient samples [18]. Calibration also manifests as a flat rank-histogram. A simple metric used below is the root-mean-squared error of the sample mean. In the large sample limit, the sample mean becomes deterministic. So we expect this error to be comparable for generative and deterministic models.
@@
The $F$ which minimizes CRPS is the true cumulative distribution of $x$. For a deterministic forecast, $F(y)=\mathbbm{1}_{\{y>=x_{0}\}}$ where $x_{0}$ is the forecast value, CRPS is equivalent to the mean absolute deviation.
@@ 6 Acknowledgements
We extend our profound appreciation to the Central Weather Administration (CWA) of Taiwan [^2] https://www.cwa.gov.tw/eng/, a premier government meteorological research and forecasting institution, for granting us access to the invaluable operational Numerical Weather Prediction (NWP) model dataset and for their expert guidance on data consultation. Our gratitude also extends to the AI-Algo team at NVIDIA, especially Kamyar Azizzadenesheli, Anima Anandkumar, Nikola Kovachki, Jean Kossaifi, and Boris Bonev for their insightful discussions. Additionally, we are indebted to David Matthew Hall, Dale Durran, Chris Bretherton for thier constructive feedback on the manuscript.
@@ 1 Our position with respect to existing works
To highlight the novel component of our work we expend our review. Table S1 above present a shortlist of the most relevant works that preform weather downscaling. Some of the successes of previous downscaling works are by preforming state vector inflation as [1], on a relatively larger domain [23], with a large resolution ratio, e.g. [24, 1] and downscaling precipitation from tropical cyclones [59]. However, most of these related works focus on a single variables per model (note that [32] provided two model each per variable and is thus listed twice in the table). The variables of interest in all these works are related to properties of cloud and precipitation. While [59] showed a successful super resolution (recovering 10km from data coarsened to 100km) of tropical cyclone precipitation. To the best of our knowledge ML downscaling of tropical cyclones, which requires accounting for different physics, across many channels and channel synthesis was not shown before.
The combined prediction of selected dynamical, thermodynamical and microphysical (cloud related) variables in concert marks a new capability of such models. It utility is demonstrated here by examining coherent structures and how all variables jointly downscaled in a physically consistent manner.
Citation
Architecture
Resolutions
Pixels
Variables
Addison et al., (2022) [1]
Diffusion
input: 60km
@@
$64^{2}$
precipitation
Harris et al., (2022) [23]
GANs +
@@
$940^{2}$
precipitation
Hatanaka et al., (2023) [24]
Cascaded
@@
Leinonen et al., (2020) [32]
GANs
input 8km
@@
$128^{2}$
precipitation
Leinonen et al., (2020) [32]
GANs
input 16km
@@
$128^{2}$
cloud optical-thickness
Price and Rasp, (2022) [43]
Corrector GAN
input 32km
@@
$128^{2}$
precipitation
Vosper et al., (2023) [59]
WGAN
input 100km
@@
Current work
CorrDiff
input 25km
@@
$448^{2}$
10 meter winds
2 meter temperature
@@
Table S1: Downscaling models presented in the most relevant works we could find with respect to the current study. We highlight the resolution ratios, the pixel size of the high resolution prediction, predicted variables and architecture.
2 Descriptions of the architecture and the channels
Table S2 lists the input and output channels for the CorrDiff model. Note that the two differ both by pixel side and by the channels themselves. For single level variables, the input includes total column water vapor but lacks the maximum radar reflectivity which is present in the output, and vice versa. The input also includes pressure level variables at the $925$, $850$, $700$ and $500$ (hpa) levels.
@@ 2 Descriptions of the architecture and the channels
Table S2: A list of the input and output resolutions and channel for the CorrDiff downscaling model. Input channels include the both single level variables and pressure level variables, the latter are used at $925$, $850$, $700$ and $500$ (hpa) levels.
Figure S1 shows the architecture of the Unet in CorrDiff.
@@ 3 Localization by two-step formulation
Figure S2 (left) demonstrates the role of the two steps associated with CorrDiff as a function of spatial scales. From the target data (blue), it is seen that the regression step learns mostly large spatial scales, leaving the small scales almost completely for the diffusion step. In addition, from Fig S2 (right), it is observed that the residual is quite localized. It seems that to fully resolve the residual at each spatial location, only 40km radius around that is needed, while for the target data there is long-range correlations up to 200km. This has important implications for training and sampling efficiency of diffusion models since one can deploy diffusion models with smaller UNet desnoing architectures to aggregate the local information. We leave further study of this for future research.
Figure S2: Left column: power spectra, right column: spatial auto-correlation. Top to bottom: maximum radar reflectivity, 10m eastward wind, 10m northward wind and 2m temperature. This figure compares the original target $x$ and the difference $r=x-\mathbbm{E}[x|y]$. The difference has greatly reduced variance at large-scales and equivalently removes the long-range auto-correlations.

### v4

@@ 5.2.1 Regression on the mean
In order to predict the conditional mean ${\boldsymbol{\mu}}=\mathbbm{E}[{\bf x}|{\bf y}]$, we employ to a UNet-regression model trained on a dataset $\{({\bf x}_{n},{\bf y}_{n})\}_{n=1}^{N}$. We utilize the specific architecture described in  [63] that incorporates attention and residual layers, allowing it to effectively capture both short and long-range dependencies in the data (see section 2 and S1). The model is optimized using a Mean-Squared-Error (MSE) loss during training.
@@ 5.3.1 Dataset
Our training and test data cover the region of Taiwan and surrounding ocean. The choice of region is driven by our partnership with a local government agency (CWA) which operates a regional dynamical downscaling. However, the region of Taiwan also presents a unique diversity of meteorological conditions and phenomena such as tropical cyclones (i.e., typhoons), mid-latitude cyclones which generate weather fronts, and steep topography with snow-capped mountains and land-sea contrast. Such diversity is hard to find elsewhere in such a relatively small domain.
@@
The input (conditioning) dataset is taken from ERA5 reanalysis [28], a global dataset at spatial resolution of about 25km and a temporal resolution of 1h, the latter matches the target data listed below. To facilitate training, we interpolate the input data onto the curvilinear grid of CWA with bilinear interpolation (with a rate of 4x), which results in $36\times 36$ pixels over the region of Taiwan. Each sample in the input dataset consists of 12 channels of information; see Table S2. This includes two pressure levels (500 hPa and 850 hPa) with four corresponding variables: temperature, eastward and northward components of the horizontal wind vector, and geopotential height. Additionally, the dataset includes single-level variables such as 2-meter Temperature, 10-meter wind vector and total column water vapor. This input channel set is admittedly somewhat arbitrary, but serves the purposes of (i) allowing a reasonable sparse representation of the thermodynamic state of the macro-scale atmosphere, while (ii) intentionally including only limited information about atmospheric water via the total vertical integral of an invisible trace gas - water vapor. This ensures the task of synthesizing radar reflectivity - a much more complex observable that relates to the sixth moment of the cloud water droplet distribution - is appropriately ambitious, as a strong test of CorrDiff’s generative component.
@@
The target dataset used in this study is a subset of the proprietary RWRF model data (Radar Data Assimilation with WRFDA  [^1] https://www.mmm.ucar.edu/models/wrfda). The RWRF model is one of the operational regional Numerical Weather Prediction (NWP) models developed by CWA [15], which focuses on radar Data Assimilation (DA) in the vicinity of Taiwan. Assimilating radar data is a common strategy used in regional weather prediction, which helps constrain especially stochastic convective processes such as mesoscale convective systems and short-lived thunderstorms. In addition, CWA assimilates several surface measurements to complement the radar data that often miss the surface values. The WRF-CWA system uses a nested 2km domain in a larger 10km domain that is driven by a global model (GFS) as boundary conditions [15]. By incorporating radar data [14], RWRF improves the short-term prediction of high-impact weather events. The radar observations possess high spatial resolution of approximately 1km and temporal resolutions of 5-10 minutes at a convective scale. These observations provide useful wind information (radial velocity) as well as hydrometeors (radar reflectivity), with a particular emphasis on the lower atmosphere. The radar data assimilation relies on the availability of reliable and precise observations, which contributes significantly to enhance the accuracy and performance of the applied deep learning algorithms in the context of NWP applications.
@@
The target dataset covers the years 2018-2021. It has a temporal frequency of one hour and a spatial resolution of 2km. We use only the inner (nested) 2km domain, which has $448\times 448$ pixels, projected using the Lambert conformal conical projection around Taiwan. The geographical extent of the dataset spans from approximately 116.371°E to 125.568°E in longitude and 19.5483°N to 27.8446°N in latitude. We sub-selected 4 channels (variables) as the target variables that are most relevant for practical forecasting: temperature at 2 meter above the surface, the horizontal winds at 10 meter above the surface and the 1h maximum derived radar reflectivity (radar reflectivity hereafter) - a surrogate of the expected precipitation. Notably, the radar reflectivity channel is not present in the input data, and needs to be predicted based on the other channels (channel synthesis). The radar reflectivity data also exhibits a distinct distribution compared to the other output channels, with positively skewed values and a prominent zero-mode consistent with typical non-raining conditions.
@@
Initially, the target data is provided in the NetCDF format, which is the output of the WRFDA assimilation process. A vertical interpolation from hybrid coordinates (i.e., sigma levels) to pressure coordinates (i.e., isobaric levels) is applied. As part of the preprocessing steps, the data is converted to the Hadoop Distributed File System (HDFS) format. Additionally, any missing or corrupted data points represented by "inf" or "nan" values are eliminated from the dataset. This leads to a reduction in the number of images from 37,944 to 33,813. See 2 for more details.
@@
To avoid over-fitting, we divide the data into training and testing sets. Three years of data 2018-2020 are used for training (24,154 images total) and 2021 is used for testing. Some selected dates from 2022 and 2023 are used for case studies.
@@ 5.3.2 Network architecture and training
The CorrDiff method has two step learning approach. To ensure compatibility and consistency, we employ the same UNet architecture for both the regression and diffusion networks. We enhance the UNet of [63] by increasing its size to include 6 encoder layers and 6 decoder layers. The base embedding size is set to 128, and it is multiplied over channels according to the multipliers [1,2,2,2,2]. The attention resolution is defined as 28. To represent time (i.e., timesteps in the diffusion process, not to be confused with the time of day), we utilize the Fourier-based position embedding. However, in the regression network, time embedding is disabled as no probability flow ODE is involved. No data augmentation techniques are employed during training. Overall, the UNet architecture comprises $80$ million parameters. Additionally, we introduce $4$ channels for sinusoidal positional embedding to improve spatial consistency, following established practices in the field [70, 19, 13].
During the training phase, we use the Adam optimizer with a learning rate of $2\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.99$. Exponential moving averages (EMA) with a rate of $\eta=0.5$ are applied, and dropout with a rate of 0.13 is utilized. We adopt the hyperparameters based on the guidelines proposed in EDM, Karras et al., (2022), including the same optimizer hyperparameters and learning rate schedule. EDM offers a physics-inspired design space based on ODEs that can be auto-tuned to our scenario (see table 1 in [33]). The only relevant hyperparameter that is tenable in this framework is $P_{mean}$ of the noise schedule. The value of this parameter was selected based on the resolution and dynamic range of the data as done in [32].
The regression network receives only the $12$ input channels from the ERA5 conditioning data. In contrast, the diffusion training concatenates these same $12$ input conditioning channels from the coarse-resolution ERA5 data with 4 noise channels to generate the output for each denoiser. To further enhance the diffusion conditioning, we also add the mean obtained by the regression model in the first stage. This addition provides more context and global information to the diffusion process, potentially improving its convergence.. For diffusion training, we adopt the Elucidated Diffusion Model (EDM), a continuous-time diffusion model. During training, EDM randomly selects the noise variance such that $\ln(\sigma(t))\sim\mathcal{N}(0,1.2^{2})$ and aims to denoise the samples per mini-batch. EDM is trained for 28 million steps, while the regression UNet is trained for only 2 million steps. The training process is distributed across 16 DGX nodes, each equipped with 8 H100 GPUs, utilizing data parallelism and a total batch size of 512. The total training time for regression and diffusion models was 7 days that amounts to approximately $21,504$ GPU-hours.
For the residual diffusion process, during training we adopt log-normal distribution for the noise standard-deviation $\sigma$; see (5) [33]. We choose $\sigma\sim{\rm lognormal}(\mu=0.0,\sigma=1.2)$ to ensure the forward diffusion completely destructs the large data intensity especially for the radar reflectivity variable. For sampling purposes, we employ the second-order stochastic sampler provided by EDM. This sampler performs $18$ steps, starting from a maximum noise variance of $\sigma_{\max}=800$ and gradually decreasing it to a minimum noise variance of $\sigma_{\min}=0.002$. We adopt the rest of hyperparamaters from EDM as listed in [33].
@@ 5.4 Evaluation criterion
Probabilistic predictions aim to maximize sharpness subject to calibration [52]. Qualitatively, calibration means that the likelihood of observing the true value is the same as observing a member drawn from the ensemble. A necessary condition for calibration is that the spread-error relationship be 1-to-1 when averaged over sufficient samples [21]. Calibration also manifests as a flat rank-histogram, both are reported in 3.4. A simple metric used below is the root-mean-squared error of the sample mean. In the large sample limit, the sample mean becomes deterministic. So we expect this error to be comparable for generative and deterministic models.
@@
here $\mathbbm{1}_{\{y\geq x\}}$ is the Heaviside step function, and $F$ which minimizes CRPS is the true cumulative distribution of $x$. For a deterministic forecast, $F(y)=\mathbbm{1}_{\{y\geq x_{0}}\}$ where $x_{0}$ is the forecast value, CRPS is equivalent to the mean absolute .
@@ 6 Acknowledgements
We extend our profound appreciation to the Central Weather Administration (CWA) of Taiwan [^2] https://www.cwa.gov.tw/eng/, a premier government meteorological research and forecasting institution, for granting us access to the invaluable operational Numerical Weather Prediction (NWP) model dataset and for their expert guidance on data consultation. Our gratitude also extends to the AI-Algo team at NVIDIA, especially Kamyar Azizzadenesheli, Anima Anandkumar, Nikola Kovachki, Jean Kossaifi, and Boris Bonev for their insightful discussions. Additionally, we are indebted to David Matthew Hall, Dale Durran, Chris Bretherton for their constructive feedback on the manuscript.
@@ 1 Our position with respect to existing works
To highlight the novel component of our work, we provide an expanded review of relevant literature. Table S1 presents a shortlist of the most relevant works that perform weather downscaling. Previous ML downscaling efforts have achieved notable successes in various areas. These include state vector inflation as [1], application to large spatial domains [26], achieving a large resolution ratio, e.g. [27, 1] and downscaling of precipitation in tropical cyclones [71]. It is worth noting, however, that the majority of these studies concentrate on downscaling a single variable per model (note that [39] provided two models, each for a single variable, and is thus listed twice in the table). The variables of interest in all these works primarily relate to cloud and precipitation properties. [71] showed a successful super resolution (recovering 10km from data coarsened to 100km) of tropical cyclone precipitation. Despite these advancements, to the best of our knowledge ML downscaling that accounts for different physics, across many channels and channel synthesis in a single model was not shown before.
The combined prediction of selected dynamical, thermodynamical and microphysical (cloud related) variables in concert marks a new capability of such models. Its utility is demonstrated here by examining coherent structures and how all variables jointly downscaled in a physically consistent manner.
Citation Architecture Resolutions Pixels Variables
Addison et al., (2022) [1] Diffusion input: 60km
@@
$64^{2}$ precipitation
Harris et al., (2022) [26] GANs +
@@
$940^{2}$ precipitation
Hatanaka et al., (2023) [27] Cascaded
@@
Leinonen et al., (2020) [39] GANs input 8km
@@
$128^{2}$ precipitation
Leinonen et al., (2020) [39] GANs input 16km
@@
$128^{2}$ cloud optical-thickness
Price and Rasp, (2022) [50] Corrector GAN input 32km
@@
$128^{2}$ precipitation
Vosper et al., (2023) [71] WGAN input 100km
@@
Current work CorrDiff input 25km
@@
$448^{2}$ 10-meter winds
2-meter temperature
@@
Table S1: Downscaling models presented in the most relevant works we could find with respect to the current study. We highlight the resolution ratios, the pixel size of the Fig. prediction, predicted variables and architecture.
2 Descriptions of the architecture and the training data
Figure S1 illustrates the architecture of the UNet, which serves a dual purpose in CorrDiff: it functions as both the regression model and the denoiser in the diffusion model. Table S2 provides a comprehensive list of the input and output channels utilized by CorrDiff. It is important to note that these input and output datasets differ not only in their pixel size but also in the specific channels they encompass. For single-level variables, the input includes total column water vapor but lacks the maximum radar reflectivity, which is present in the output, and vice versa. The input also includes pressure level variables at the $850$ and $500$ (hPa) levels, which combine to 12 input channels.
@@ 2 Descriptions of the architecture and the training data
Table S2: A list of the input and output resolutions and channel for the CorrDiff downscaling model. Input channels include the both single level variables and pressure level variables, the latter are used at $850$ and $500$ (hPa) levels.
Figure S2 below shows a time series of the normalized target data, where the mean and the 2 and 4 standard deviations from the mean are plotted. A single corrupted data point is evident by the negative spike in 2m temperature. Some missing data periods are also evident by gaps, such as around May 2021; these were removed due to ’NaNs’ or ’Inf’ in either input or target data. The validation data constitutes the last 12 months from 2021-01-01 00:00:00 UTC. Moreover, the training data (2018-2020) has 22 hourly snapshots from 5 named typhoons of category 1 or higher in the domain. As a rough estimate, frontal systems seem to pass through the domain between 4-8 time per year (every 2-3 weeks in the winter) and typically last about 8h within the domain. Therefore, both typhoons and frontal systems are rather rare in the training data.
Figure S2: Time series of the mean (black line), mean ±$\pm 2$ standard deviation (red shading) and mean ±$\pm 4$ standard deviation (grey shading) for the four target channels as used by the model. This analysis is after re-scaling the channels by subtracting their global mean and dividing by global standard deviation. Statistics is computed over the domain at each date and time in the dataset.
@@ 3 Localization by two-step formulation
Figure S3 (left) demonstrates the role of the two steps associated with CorrDiff as a function of spatial scales. From the target data (blue), it is seen that the regression step learns larger spatial scales, leaving some of the small scales for the diffusion step. In addition, from Fig. S3 (right), it is observed that the residual is more localized and varies less overall, especially for the temperature field, which is strongly driven by changes in elevation. This has important implications for training and sampling efficiency of diffusion models as one can deploy diffusion models, with smaller UNet denoising architectures to aggregate the local information. We leave further study of this for future research.
Figure S3: Left column: power spectra, right column: spatial auto-correlation. Top to bottom: maximum radar reflectivity, 10m eastward wind, 10m northward wind and 2m temperature. This figure compares the original target $x$ and the difference $r=x-\mathbbm{E}[x|y]$. The difference has reduced variance at large-scales and equivalently removes the long-range auto-correlations.